OpenStudy (anonymous):

A long distance trucker traveled 160 miles in one direction during a snow storm. The return trip in rainy weather was accomplished at double the speed and took 4 hours less gime. The going speed was ?? mph.

OpenStudy (anonymous):

20 mph

OpenStudy (anonymous):

lets see why. time is \[\text {time}=\frac{\text{distance}}{\text{rate}}\] the distance both directions is 160. if the rate going it r, then the rate returning is 2r (double) and you are told that the time returning is 4 hours less than the time going. so an equation could be \[\frac{160}{r}=\frac{160}{2r}+4\]

OpenStudy (anonymous):

solve this for r to get 20 and the check is easy. going took \[\frac{160}{20}=8\] and \[\frac{160}{40}=4\]

OpenStudy (anonymous):

Thank you!